# Beyond Sequence: Impact of Geometric Context for RNA Property Prediction

### v1

@@ 2.3 Models
    Transformer1D: The Transformer1D model is a standard Transformer architecture adopted for RNA sequence processing. It includes an embedding layer to convert input tokens into dense vectors, positional encoding (PE) to retain sequence order, and a multi-layer Transformer encoder to capture long-range dependencies within the sequence.
@@
    SchNet: An SE(3)-invariant network designed for molecular property prediction on geometric graphs of atomic systems. It operates by modeling interactions through continuous-filter convolutional layers. Since the continuous-filter in Schnet is conditioned on distance features, it maintains invariance to rotations and translations of atom coordinates.
@@
    E(n)-Equivariant Graph Neural Network (EGNN): An equivariant network for geometric graphs with rotations, translations, and reflections symmetry. The EGNN operates by as a non-linear message passing between scalar-invariant and vector equivariant quantities.
Training and evaluation
All models were trained on a NVIDIA A100 GPU. To ensure hyperparameter parity for each baseline, hyperparameters were optimized using Optuna [1], restricting the search to models with fewer than 10 million parameters that fit within the GPU memory constraint of 80GB. All model hyperparameters, training, and evaluation details are reported in Appendix E. We ran all models for $5$ random data splits (train:val:test split of 70:15:15) and we report average performance with a standard deviation across splits. The mean column-wise root mean squared error (MCRMSE), introduced in [58], is used as the evaluation metric. It is defined as $\text{MCRMSE}(f,D)=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(\hat{y}_{i}-y_{i}\right)^{2}}$, where $f$ represents a model, $D$ is the dataset, and $\hat{y}_{i}$ and $y_{i}$ are the predicted and true values for data point $i$.
3 Task Definitions
Task 1: Impact of structural information on prediction performance
Task 2: Model efficiency in limited training data settings
Acquiring high-quality comprehensive RNA datasets is often challenging and resource-intensive thus limiting the amount of labeled data for training [51, 7]. This task aims to investigate how model performance depends on the amount of training data used, evaluating the sample efficiency of each family of models. In other words, given a dataset $D=\{X,Y\}$, let $D_{\alpha}=\{X_{\alpha},Y_{\alpha}\}$ be a subset where the training set is reduced to a fraction $\alpha$. We train models on different sets of $D_{\alpha}$ datasets with decreasing $\alpha$.
Task 3: Performance with partial sequence labeling
Due to the high cost of measuring properties for every nucleotide in RNA sequence, real-world datasets often contain partial annotations [60] where labels are only available for the first small part of the sequence. This task is relevant for nucleotide-level datasets and it aims to investigate how well a model can generalize to a whole RNA sequence when labels are only available for a portion of it.
Task 4: Robustness to sequencing noise
Acquiring RNA data requires sequencing. In practice sequencing procedures may introduce sequencing errors (random mutations of nucleotides) that vary depending on the sequencing technology and platform [33, 14]. These errors affect the raw sequence data and propagate to structural noise in 2D and 3D. The goal of this task is to assess how well models can maintain reliable performance when trained and tested under the same distribution of realistic levels of sequencing noise observed in practice, ensuring robustness across a consistent noise environment. This reflects real-world cases where a specific sequencing method produces noisy data, but the noise characteristics are stable across training and deployment.
Task 5: Generalization to Out-of-Distribution (OOD) data
This task focuses on a different practical challenge: models trained on high-quality RNA sequences are often deployed in conditions where the data exhibits different noise characteristics due to batch effect [53] or the use of different sequencing platforms [52]. Here, the objective is to evaluate how well models generalize to OOD datasets with different levels of sequencing noise, assessing the extent of performance degradation as noise levels increase. This task simulates the scenario where a model encounters noisier data than it was trained on, highlighting its ability to adapt to unexpected experimental conditions.
4 Experiments and Results
4.1 Impact of explicit geometry learning on model performance
We begin by addressing Task 1, where we compare the performance of model families when trained and evaluated on the downstream RNA datasets. Additionally, we provide runtime and memory comparison in Appendix B.
Model COVID Ribonanza Tc-Ribo Fungal
1D model
2D model
Transformer1D2D 0.305$\pm$0.012 0.514$\pm$0.004 0.633$\pm$0.001 OOM
GCN 0.359$\pm$0.009 0.595$\pm$0.006 0.701$\pm$0.004 1.192$\pm$0.077
GAT 0.315$\pm$0.006 0.534$\pm$0.006 0.685$\pm$0.024 1.112$\pm$0.035
ChebNet 0.279$\pm$0.007 0.468$\pm$0.002 0.621$\pm$0.022 0.973$\pm$0.003
3D model
EGNN (w/o pooling) 0.480$\pm$0.025 0.808$\pm$0.023 0.725$\pm$0.002 OOM
SchNet (w/o pooling) 0.499$\pm$0.003 0.843$\pm$0.004 0.696$\pm$0.008 OOM
EGNN (nuc. pooling) 0.364$\pm$0.003 0.619$\pm$0.007 0.663$\pm$0.010 OOM
SchNet (nuc. pooling) 0.390$\pm$0.006 0.685$\pm$0.006 0.655$\pm$0.038 OOM
2D models consistently outperform 1D model
Results in Table 1 reveal that 2D methods consistently outperform the 1D sequence model across all datasets. Notably, the Transformer1D2D model, which simply augments the attention matrix with adjacency features alongside, achieves around 10% lower prediction MCRMSE on average across datasets than its geometry-free counterpart. This suggests that explicitly incorporating structural information is crucial, as learning from sequence data alone proves to be insufficient. Further experiments, detailed in the Appendix C.4, investigate the learned attention maps of both the Transformer1D2D and the Transformer1D model and their correlation with structural information and reveal that Transformer1D2D attention maps are much more closely aligned with the topological structure of nucleotide graph, reinforcing the conclusion that explicit encoding of structural information is essential for improved performance.
Spectral GNN outperforms spatial GNNs in 2D
ChebNet, a spectral method, outperforms spatial methods such as GCN, GAT, Graph Transformer, and GraphGPS, achieving a prediction MCRMSE 2.5% lower than the next best 2D model across datasets. Spatial GNNs aggregate node features layer by layer, emphasizing local information within a fixed distance. While computationally efficient, these methods are limited by the 1-Weisfeiler-Lehman (WL) test, which constrains the expressive power of node-based updates [63]. In addition, spatial GNNs may suffer from a limited receptive field while spectral methods approximate global graph features, enabling a global receptive field since the first layer. This allows ChebNet to effectively process global information which is important for RNA data due to potential long-range interaction between nucleotides.
Challenges of modeling geometric context in all-atom resolution
Contrary to our expectations, 3D models at all-atom resolution (EGNN w/o pooling and SchNet w/o pooling in Table 1) show relatively high prediction MCRMSE across datasets, underperforming compared to 1D and 2D methods. We hypothesize this is due to two factors.
First, all-atom SchNET and EGNN rely on a limited local neighborhood of adjacent atoms, limiting their receptive fields and preventing them from capturing long-range dependencies (see Appendix Table C.1), which can be crucial for determining RNA properties [46, 3]. Expanding the local neighborhood for these methods turns challenging due to the overwhelming scale of large molecular systems in all-atom resolution. Second, the performance of 3D models is often limited by the inherent inaccuracies in 3D structure prediction tools, which are generally less reliable compared to 2D structure prediction methods [37].
To address the receptive field limitation of all-atom methods, we employ a biological prior by pooling atomic features into nucleotide-level representations after a few layers of all-atom operations. This strategy is aligned with RNA’s natural secondary structure, where atoms group into nucleotides, and nucleotides form the complete RNA molecule [12]. This novel strategy allows us to maintain all-atom resolution in the initial layers while increasing the receptive field via pooling. By balancing the number of all-atom and nucleotide layers as hyperparameters, we can balance fine- and coarse-grained all-atom and nucleotide resolution. We call 3D models with pooling EGNN (nuc. pooling) and SchNet (nuc. pooling). This strategy significantly enhances 3D model performance, reducing prediction MCRMSE by $\sim$10compared to the original EGNN and SchNet, and outperforming 1D models on the Ribonanza-2k and Tc-Riboswitches datasets by 5% on average. On the COVID dataset, EGNN (nuc. pooling) matches the Transformer1D model but still trails behind 2D models. All subsequent experiments report results using nucleotide-pooled versions of the 3D models.
Next, we investigated the second hypothesis regarding higher noise in 3D structures by quantifying variability in predicted structures across different 3D prediction tools in Appendix D. We observe substantial variability (between 11-45Å RMSD across structures given by different 3D structure prediction tools), suggesting considerable noise in 3D predictions, which likely contributes to the poorer performance of 3D models.
We also evaluate the impact of partial property labels for nucleotide-level tasks, a common occurrence owing to costly experimental measurements [58] to identify which models are best suited to handle the challenges of incomplete labels in RNA property prediction. For this, we use the COVID and Ribonanza datasets as these datasets contain nucleotide-level labels. We train the models using all training data but with varying proportions of labeled nucleotides (20%, 40%, 60%, 80%, and 100%) per sequence, thus simulating incomplete or sparse labeling, while testing on fully labeled test sets (Appendix Fig. 8(b) for illustration).
(a) COVID
(b) Ribonanza
(b) Ribonanza
Figure 2: Performance versus fraction of training data across various datasets. Model performance improves with increasing data, with lower MCRMSE across all models. 2D models consistently outperform 1D models, particularly in low-data regimes, underscoring the value of structural information for generalization. Dotted, solid, and dashed lines denote 1D, 2D, and 3D methods, respectively, which apply consistently throughout all figures in this paper.
(a) COVID
(b) Ribonanza
More training data improves performance
Unsurprisingly, across all models and datasets, a clear trend emerges: increasing the amount of available data, whether through higher number of training data points or greater proportion of available labels leads to improved performance (Fig. 2, Fig. 3, Appendix Tables in Sec. F.1F.2). However, the degree of performance improvement varies significantly between model types, as analyzed next.
2D models excel in low data and partial label regimes
Noise Ratio
0 0.1 0.2 0.3
1D Refer to caption Refer to caption Refer to caption Refer to caption
2D Refer to caption Refer to caption Refer to caption Refer to caption
3D Refer to caption Refer to caption Refer to caption Refer to caption
4.3 Model Robustness and Generalization Under Sequencing Noise
Tasks 4 and 5 both deal with model performance with noise in the data, but focus on different aspects, robustness to noise, and ability to generalize across unseen noise distributions. As explained in Sec. 3, sequencing noise is common depending on the sequencing method and platform used [14], thereby introducing errors in sequences that propagate into 2D and 3D structures. Additionally one of the deployment scenarios involves models trained on high-quality clean data applied for datasets acquired under noisy conditions owing to different sequencing platforms or experimental batch effects [52].
  • Robustness: We introduce sequencing noise into the training, validation, and test sequences to simulate realistic sequencing errors. Nucleotide mutations are applied with probabilities {0.05, 0.1, 0.15, 0.2, 0.25, 0.3}, mirroring typical sequencing error rates [36] which also propagates to the 2D and 3D structures (Fig. 4) via structure prediction tools. Importantly, these mutations are not random; the likelihood of a particular nucleotide mutating into another varies, as is well documented in sequencing studies [36]. Our noise model reflects these real-world mutation profiles. Crucially, while the input training, validation, and test sequences contain noise, the property labels remain clean. This again reflects practical scenarios where labels are experimentally determined independent of sequencing and thus unaffected by sequencing errors.
(a) COVID
(b) Ribonanza
(c) Fungal
(d) Tc-Riboswitches
(a) COVID
(b) Ribonanza
(c) Fungal
(d) Tc-Riboswitches
Expectedly, across both tasks, increased noise levels generally lead to worse test MCRMSE for all models, indicating a decline in prediction performance (Fig. 5, Fig. 6 and Appendix Tables in Sec. F.3F.4).
Among all models, Transformer1D demonstrates the highest robustness and generalization, exhibiting the least performance degradation as noise levels increase. Notably, in generalization experiments, Transformer1D achieves the best prediction MCRMSE on the COVID, Ribonanza, and Tc-Riboswitches datasets under higher noise levels (Fig. 6 and Appendix F.4). The reliance of Transformer1D on sequence-only information without considering geometric context, while being a weakness in other scenarios, becomes a strength in case of noisy sequences as minor sequencing errors may severely alter the predicted RNA structures (Fig. 4). While its performance on the Fungal dataset is worse overall, Transformer1D maintains remarkably consistent performance as noise increases. Transformer1D2D ranks just behind Transformer1D outperforming other 2D and 3D models. This can be attributed to Transformer1D2D’s ability to selectively focus on sequence information for noisy data rather than relying on structural data alone as the self-attention is only weakly conditioned on the graph topology.
More elaborated 2D and 3D models, which rigidly rely on structural information, are significantly more affected by noise, underperforming compared to plain sequence baseline in both robustness and generalization experiments. Interestingly, ChebNet shows the worst generalization among 2D models, although it performs on par with other methods in robustness experiments. This suggests that while ChebNet struggles with OOD noise, its performance gets better when it is also trained on the same noise level used during testing, highlighting the need for retraining for different experimental data batches/noise levels in real-world applications. Across both experiments, 3D models have poor performance for more noisy conditions, particularly with the COVID and Ribonanza datasets, due to their dependence on 3D structures which are also sensitive to the propagation of sequencing errors (Fig.4 bottom). In the smaller Tc-Riboswitches dataset, model performances vary more, likely due to limited data size, but transformer models still consistently demonstrate greater robustness to noise.
5 Conclusion
We present the first comprehensive study on the benefits and challenges of the effect of geometric context for RNA property prediction models. By providing a curated set of RNA datasets with annotated 2D and 3D structures, we systematically evaluate the performance of 1D, 2D, and 3D models under various real-world conditions, such as limited data, partial labeling, sequencing errors, and out-of-distribution generalization. Our results reveal that 2D models outperform 1D and 3D models, with spectral graph neural networks excelling even in low-data and partial labeling scenarios. For 3D models, we find that their potential benefits are hindered by the limited receptive field, computational complexity, and structural noise from RNA structure prediction tools. At the same time, 1D models demonstrate better robustness compared to 2D and 3D models in noisy and OOD conditions. This study highlights the value and limitations of using geometric context for RNA modeling. Future work could focus on ensembling 1D, 2D, and 3D models for complementary strengths, and on improving 2D and 3D models to better handle noise from structure prediction tools.
Appendix Overview
  • Appendix A: Related Work.
  • Appendix B: Memory and Computational Constraints.
  • Appendix C: Additional Experimental Information.
  • Appendix D: Analysis of Noise in 3D Structures.
  • Appendix E: Reproduction.
  • Appendix F: Additional Results.
Appendix A Related Work
RNA property prediction
RNA-specific models remain scarce, likely due to limited specialized datasets. Recent advancements in sequence modeling have shown promise, particularly with foundation models like RNA-FM [8], UTRBERT [66], and RINALMO [35], which use transformer-based architectures pre-trained on large RNA sequence corpora to predict various RNA functions and structures. While RNNs and CNNs have been applied to tasks like RNA methylation and protein binding [56], they struggle with long-range dependencies. Hybrid models like RNAdegformer [21], combining convolutional layers with self-attention, improve predictions by capturing both local and global dependencies. Although some efforts integrate 2D structures with transformers, explicit 2D and 3D geometric modeling for RNA remains underexplored, with graph-based models mainly focusing on RNA-protein and RNA-drug interaction tasks rather than property prediction [29, 65, 4, 69].
RNA structure prediction
RNA 2D structure prediction has progressed from dynamic programming methods like Vienna RNAfold [23] to deep learning-based tools like SPOT-RNA2 [48] and UFold [16], which enhance accuracy by using neural networks and evolutionary data. Models such as E2Efold [9] and RNA-FM [8] employ transformer architectures to achieve state-of-the-art results in secondary structure prediction.
RNA 3D structure prediction has progressed through ab initio, template-based, and deep learning approaches. Ab initio methods (e.g., iFoldRNA [43], SimRNA [6]) balance detail and efficiency but struggle with non-canonical interactions. Template-based models (e.g., FARNA/FARFAR [10], 3dRNA [68]) depend on existing structures but are limited by available data. Deep learning models like DeepFoldRNA [34], RhoFold [45], RoseTTAFoldNA [5], and trRosettaRNA [57] show promise in predicting 3D structures from sequence data but face challenges with novel RNA families due to RNA’s conformational flexibility [30].
Appendix B Memory and Computational Constraints
@@
Appendix C Additional Experimental Information
C.1 Dataset Statistics
@@
C.2 Comparison of partial training data and partial sequence labeling
@@
C.3 Details of noisy experiments: robustness and generalization
@@
C.4 Analysis of Transformer1D and Transformer1D2D

### v2

@@ Appendix B Models overview
    Transformer1D: The Transformer1D model is a standard Transformer architecture adopted for RNA sequence processing. It includes an embedding layer to convert input tokens into dense vectors, positional encoding (PE) to retain sequence order and a multi-layer Transformer encoder to capture long-range dependencies within the sequence.
@@
    SchNet: An SE(3)-invariant network designed for molecular property prediction on geometric graphs of atomic systems. It operates by modeling interactions through continuous-filter convolutional layers. Since the continuous filter in Schnet is conditioned on distance features, it maintains invariance to rotations and translations of atom coordinates.
@@
    E(n)-Equivariant Graph Neural Network (EGNN): An equivariant network for geometric graphs with rotations, translations, and reflections symmetry. The EGNN operates as a non-linear message passing between scalar-invariant and vector equivariant quantities.
  10. 10.
    FAENet: FAENet, or Frame Averaging Equivariant Network, is a lightweight and scalable Graph Neural Network designed for materials modeling. It introduces Stochastic Frame Averaging to achieve E(3)-equivariance through data transformations rather than architectural constraints, enabling superior accuracy and scalability on tasks like molecular and materials property prediction.
  11. 11.
    GVPGNN: Geometric Vector Perceptron GNN is a graph-based model designed to learn from 3D molecular structures by integrating geometric and relational information. It uses Geometric Vector Perceptrons to encode scalar and vector features, enabling effective modeling of both protein structure and interactions.
  12. 12.
    DimeNet: Directional Message Passing Neural Network leverages directional message passing to encode both distance and angular information between atoms. It utilizes spherical harmonics and Bessel functions for rotation equivariant message representations, enabling precise modeling of molecular interactions.
  13. 13.
    FastEGNN: FastEGNN introduces virtual nodes to traditional equivariant GNNs, enabling global message passing while maintaining scalability. These virtual nodes are designed to approximate and enhance real node interactions, ensuring E(3)-equivariance for efficient modeling of large geometric graphs.
Appendix C Memory and Computational Constraints
@@
Appendix D Detailed Task Descriptions
To comprehensively evaluate RNA property prediction models, we define five tasks that simulate real-world experimental challenges. These tasks are designed to assess the performance, efficiency, and robustness of models under various conditions encountered in RNA research and application. Below, we provide detailed explanations for each task, their scientific motivation, and their relevance to practical RNA property prediction.
D.1 Task 1: Impact of Structural Information on Prediction Performance
Motivation: RNA molecules fold into complex secondary (2D) and tertiary (3D) structures that dictate their functional properties. While traditional ML approaches often rely solely on the nucleotide sequence (1D representation), it is well-known in chemistry and biology that structural information influences RNA properties (Schlick & Pyle, 2017). However, the extent to which geometric representations can improve performance of ML models remains unclear.
Task Definition: We evaluate multiple models across three classes (models described in main text):
  • 1D models: Utilize only the linear sequence of nucleotides.
  • 2D models: Incorporate RNA secondary structure, such as base-pairing interactions.
  • 3D models: Leverage full tertiary structural information, including spatial arrangements of nucleotides.
By comparing their performance across diverse datasets, this task quantifies the benefits of using geometric information. Specifically, we aim to determine if models with explicit 2D or 3D inputs outperform sequence-only models, thereby highlighting the added value of structural representations in RNA property prediction.
Significance: This analysis addresses the gap in existing literature, where the impact of structural information on RNA property prediction remains underexplored. The findings inform whether additional computational costs associated with structural modeling are justified by performance gains.
D.2 Task 2: Model Efficiency in Limited Training Data Settings
Motivation: High-quality RNA datasets with experimentally measured properties are scarce due to the technical difficulty and cost of experimental data acquisition (Byron et al., 2016; Teufel & Sobetzko, 2022). In such scenarios, models must be sample-efficient, learning effectively from small datasets.
Task Definition: Let $D=\{X,Y\}$ represent a dataset, where $X$ is the RNA input and $Y$ is the property label. For this task, we define subsets $D_{\alpha}$ by sampling a fraction $\alpha$ of the original dataset. We evaluate models trained on varying $\alpha$ values (e.g., $\alpha=25\%,50\%,75\%$) to assess how performance scales with reduced training data.
Significance: This task simulates real-world scenarios where large labeled datasets are unavailable. It evaluates the ability of different models to generalize from limited data, providing insights into their efficiency and practical applicability.
D.3 Task 3: Performance with Partial Sequence Labeling
Motivation: In nucleotide-level RNA datasets, labels for molecular properties (e.g., binding affinities or reactivity) are often available only for a subset of nucleotides in a sequence (Wayment-Steele et al., 2022b). This is due to the high cost of annotating every nucleotide experimentally.
Task Definition: We consider datasets where only the first $\ell$-nucleotide positions in each sequence are labeled, representing partial sequence labeling. Models are trained using these incomplete labels and evaluated on their ability to predict properties across the full sequence. Performance is assessed using datasets with varying $\ell$-label fractions.
Significance: This task reflects a critical real-world challenge where complete annotations are unavailable. It assesses a model’s ability to generalize from partial labels to unseen portions of RNA sequences, a valuable capability for applications with sparse experimental data.
D.4 Task 4: Robustness to Sequencing Noise
Motivation: RNA sequencing technologies often introduce noise in the form of random nucleotide errors (e.g., insertions, deletions, or substitutions) (Ozsolak & Milos, 2011; Fox et al., 2014). This noise propagates through derived RNA structures (2D and 3D), potentially degrading model performance. Understanding how well models handle noise is crucial for their reliable deployment.
Task Definition: To simulate realistic sequencing errors, we introduce controlled noise levels into RNA sequences during both training and testing. Models are evaluated for their ability to maintain performance under consistent noise conditions, representing a scenario where noise characteristics are stable across experimental phases.
Significance: This task assesses a model’s robustness in practical settings where sequencing noise is unavoidable. It provides insights into the resilience of RNA models to variations in input quality.
D.5 Task 5: Generalization to Out-of-Distribution (OOD) Data
Motivation: RNA models are often trained on high-quality experimental datasets but deployed in conditions where data characteristics differ significantly due to variations in sequencing platforms or protocols (Tran et al., 2020; Tom et al., 2017). This mismatch can lead to performance degradation.
Task Definition: We evaluate models trained on clean RNA data and tested on datasets with higher levels of noise (representing OOD conditions). Performance metrics are analyzed as a function of increasing noise, quantifying the models’ ability to generalize to unseen distributions.
Significance: This task mirrors real-world deployment scenarios where models encounter noisy or biased data. It highlights the limitations of models trained on idealized datasets and informs strategies for improving generalization under OOD conditions.
Appendix E Discussion on Advanced Model Architectures
E.1 Enhancing 3D RNA Prediction with High-Degree Steerable Features
Apart from the 3D models considered in this work, models using high-degree steerable features such as TFN (Thomas et al., 2018), MACE (Batatia et al., 2022) and NequIP (Batzner et al., 2022) represent an important aspect of equivariant models, with potential to enhance model expressiveness by incorporating higher-order information. While memory constraints have limited their practical implementation in our current work, techniques such as scalarization of high-degree steerable features, as demonstrated in works such as HEGNN (Frank et al., 2024) and SO3krates (Cen et al., 2024), could address this challenge. For instance, HEGNN can improve the expressivity of equivariant GNNs by mitigating expressivity degeneration on symmetric graphs and leveraging higher-order representations. Although 3D RNA graphs are too complex to be considered symmetric, the ability to handle high-degree steerable features in large RNA graphs, with hundreds of nucleotides and thousands of atoms, remains valuable and has the potential to improve performance. Future work can explore such methods to improve 3D RNA prediction.
E.2 Mitigating Sequencing Noise
Notably, most studies in this field do not explicitly address sequencing noise common for RNA in model architecture design, prompting a need to explore effective strategies for mitigating its impact. An effective way to handle noise can be through ensemble methods. For instance, combining 1D, 2D, and 3D models by independently learning representations and integrating them via attention mechanisms that dynamically weigh each modality based on task relevance and noise can leverage the robustness of 1D methods while benefiting from the strengths of 2D and 3D approaches and be a good direction for future research.
Appendix F Additional Experimental Information
F.1 Dataset Statistics
@@
Task Level RNA-level Nucleotide-level Nucleotide-level RNA-level
@@
F.2 Comparison of partial training data and partial sequence labeling
@@
F.3 Details of noisy experiments: robustness and generalization
@@
F.4 Analysis of Transformer1D and Transformer1D2D